Solve equations by factoring
Solve the equation.
3x^2 + 9x = 0

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
3x^2 + 9x = 0
3x(x+3)=0
3x = 0
x = 0
(x+3)=0
x = -3